40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Feb 19 13:28:44 2010, Koji, Update, General, Prep for Power Supply Stop  screen_shot.png
    Reply  Sun Feb 21 17:44:35 2010, rana, Update, General, Prep for Power Supply Stop  
       Reply  Mon Feb 22 07:25:58 2010, rana, Update, General, Prep for Power Supply Stop  
          Reply  Mon Feb 22 09:45:34 2010, josephb, Update, General, Prep for Power Supply Stop  
             Reply  Mon Feb 22 11:42:48 2010, Koji, Update, General, Prep for Power Supply Stop  
Message ID: 2620     Entry time: Sun Feb 21 17:44:35 2010     In reply to: 2617     Reply to this: 2621
Author: rana 
Type: Update 
Category: General 
Subject: Prep for Power Supply Stop  
  • Turned on the RAID attached to linux1 (its our /cvs/cds disk)
  • Turned on linux1 (it needed a keyboard and monitor in order to be happy - no fsck required)
  • Turned on nodus (and started ELOG) + all the control room machines
  • Turned on B/W monitors
  • Untaped fridge


  • Found several things OFF which were not listed in the Wiki...
  • Turned ON the 2 big isolation transformers (next to Steve's desk and under the printer). These supply all of the CDS racks inside.
  • ~75% of the power strips were OFF in the CDS racks ?? I turned on as many as I could find (except the OMC).
  • Switched on and keyed on all of the FE and SLOW crates in no particular order. Some of the fans sound bad, but otherwise OK.
  • Turned on all of the Sorensens that are labeled.
  • Turned ON the linear supplies close to the LSC rack.
  • ON the Marconis - set them according to the labels on them (probably out-dated).
  • After restoring power to the PSL enclosure (via the Isolation Transformer under the printer) turned the Variac ON and HEPA on full speed.
  • Plugged in the PSs for the video quads. Restored the Video MUX settings - looks like we forgot to save the correct settings for this guy...


PSL


1) Turned on the chiller, then the MOPA, then the RC's Heater power supply.
2) Shutter is open, laser is lasing, PMC is locked.
3) RC temperature is slowly rising. Will probably be thermalized by tomorrow.

Sun Feb 21 20:04:17 2010
Framebuilder is not mounting its RAID frames - in fact, it doesn't mount anything because the mountall command is failing on the RAID with the frames. The Jetstor RAID is also not responding to ping. Looks like the JetStor RAID which has all of our frames is still on the old 131 network, Joe.
ELOG V3.1.3-